# Multi-task Planar Reconstruction with Feature Warping Guidance

### v1

@@ 4.1 Setup & Training details
For comparison between different model versions, we train a base model initialized with a pretrained ResNet-50 backbone and employ a data augmentation scheme where each sample has a 15% chance of undergoing one of several augmentations, such as a) jitter of brightness, contrast, hue, saturation, b) Planckian jitter (Zini et al.,, 2023), c) Gaussian noise, or d) motion blur. We use learning rate warm-up for the first 2000 steps starting from a learning rate of $1\mathrm{e}{-6}$ and increases until $2\mathrm{e}{-4}$. After the initial warm-up period, the learning rate is reduced by a factor of 0.1 given no improvement to the validation loss. For quicker and more fair comparison of model variations, a base model with the best validation loss was saved at epoch 9 and used as initialization to our main models, which were trained for an additional 11 epochs each. We employ early stopping if validation loss fails to improve for 5 consecutive epochs and save the model with best validation performance as well as the last checkpoint. For evaluation, we take the best of either saved model. The additional models trained using the base model initialization do not use data augmentation, and has 500 steps of learning rate warm-up starting from $1\mathrm{e}{-6}$ to $1\mathrm{e}{-5}$. We use a batch size of 32 for the single view model with gradient accumulation to mitigate the higher instability associated with multi-task learning. We train the models on a single NVIDIA Ampere A100 GPU. For evaluation and FPS calculation, we use a single NVIDIA GeForce RTX 3090 GPU for all models.
Table 1: Model comparison results on ScanNet dataset for variations of the SoloPlane model and other single-image planar reconstruction methods.
@@
PlaneRCNN 0.165 0.07 0.278 0.187 0.78 0.954 0.991 0.193 - 7
Solop-5lvls 0.143 0.059 0.276 0.185 0.813 0.96 0.99 0.416 0.314 38
Solop-SV 0.134 0.052 0.259 0.178 0.832 0.964 0.991 0.389 0.267 43
Solop-MV 0.136 0.054 0.261 0.177 0.832 0.962 0.991 0.427 0.344 43
Solop-MV-gw 0.133 0.052 0.259 0.177 0.833 0.964 0.992 0.434 0.347 43
@@ 4.2 Dataset
We use the ScanNet dataset which contains RGB-D images from video sequences, containing 2.5 million views complete with camera parameters and instance level semantics (Dai et al.,, 2017). The ground truth plane instance annotations for instance masks and plane parameters are generated by the authors of PlaneRCNN, and we follow the same process for filtering and preprocessing the planes (Liu et al., 2018a, ). We also obtain the corresponding plane instance semantics from the metadata of the plane annotations. The ground truth plane data often exhibited issues such as over-segmented, rough edges, or missing plane instances, as planes with a depth error above a $0.1$ meter threshold were omitted. For multi-view guidance training, we take sample pairs which are 10 time-steps away. In some cases, a neighbouring ground truth plane image might contain a segment which is missing in the source view, and vice versa. For the single-view model, we use 60,000 random samples from the training set and 10,000 from the validation set. For the multi-view model, we use 30,000 neighboring pairs for training and 5,000 pairs for validation.
@@ 4.3 Comparison
Our model is most directly comparable to the PlaneAE model by Yu et al., (2019), primarily due to the speed of prediction and the fact that they predict plane parameters directly using a single image as input. Given the inconsistent quality of ground truth plane data, the authors of PlaneMVS manually selected stereo pairs for the test set, which contained samples with more complete plane annotations (Liu et al.,, 2022). We run our evaluations using the same test set. For a fair comparison, we train the PlaneAE model for a total of 20 epochs using a ResNet-50 backbone and the same data with an input size of 480 x 640. The original model was trained using an input size of 192 x 256, resulting in a higher FPS. To align with our training regimen, we train PlaneAE for 11 epochs using 60,000 samples and an additional 9 epochs with 100,000 random samples. We retain the original training configuration of the authors (Yu et al.,, 2019). While PlaneRCNN also uses a form of multi-view warping loss during training and takes a single image at inference time, its slower inference speed makes it a less direct comparison. We run evaluations on the provided model from authors Liu et al., 2018a .
Figure 5: Per-pixel recall at varying depth thresholds in meters.
@@ 4.4 Evaluation Metrics
We follow previous methods (Yu et al.,, 2019; Liu et al., 2018a, ) and calculate the per-pixel depth recall at varying thresholds in meters, shown in Fig. 5. We also calculate standard depth and detection metrics for a comprehensive evaluation of model performance. Average Precision (AP) is used to assess the quality of the predicted masks, and Mean Average Precision (mAP) takes into account the semantic labels by averaging AP across class categories. We also calculate model efficiency using Frames Per Second (FPS). The results of these evaluations are summarized in Table 1, which shows a marked improvement using our architecture.
@@ 4.5 Results
The task of segmentation becomes more challenging when predicting multiple classes, as overlapping masks from different classes are less likely to be suppressed. The oversegmentation issue seemed to be more pronounced in the single view model, whereas multi-view guidance using plane features helped to produce more complete and less over-segmented masks. This improvement is likely attributable to feature sharing and the correlation between ground truth plane instance masks and plane parameters.
@@
Quantitative All results are obtained using the selected test set chosen by the authors of PlaneMVS. The authors Liu et al., (2022) manually selected a higher quality set to evaluate on due to the incomplete and imprecise nature of the ground truth plane annotations. The resulting test set contains 949 image pairs. Our quantitative findings from model comparisons, summarized in Table 1, indicate that our multi-view model variant not only matches the performance of the single-view model in depth metrics, but also shows a significant improvement in detection metrics. This demonstrates the efficacy and improved data efficiency in using multi-view guidance via warping in feature space, at least in the case of using shared features for multitask learning. Since all SoloPlane variants use a single image at inference time, the FPS result is the same for the versions of the model using 3 feature levels (Solop-SV, Solop-MV, Solop-MV-gw), but significantly reduced for the version with the original 5 level architecture (Solop-5lvls). Solop-MV-gw achieves better depth recall comparatively (see Fig. 5), while all Solop variants outperform the comparison models on standard metrics.
Qualitative In contrast to previous works that predicted a binary plane indication, the incorporation of multi-class semantics introduces an added complexity. The change made to the focal loss for category predictions (see Section 3.2) leads to more confident scoring as well as a potential increase in false positives, which is already exacerbated in the case of multi-class predictions. However, we found that raising the score threshold for the final masks partially mitigated this issue. See Fig. 4 for visual results. The structure of the scene is easier to predict than the exact depth, a challenge presented when using a single image for inference. The visible differences between predictions and ground truth result mainly from depth errors, as normal errors occur largely around the edges (see Fig. 7 for a 3D visualization of depth and surface normal errors). Sample visualizations of the semantic predictions can be found in Fig. 6. Cases of oversegmentation can occur due to prediction of different classes, or different plane orientation, as each mask represents a planar segment associated with a class label. Overall, our model demonstrates robust performance both visually and quantitatively for the task of planar reconstruction with semantic labels.
@@ 5 DISCUSSION AND FUTURE WORK
In this work, we introduce SoloPlanes, a real-time semantic planar reconstruction network which shows cross-task improvement when using multi-view guidance in feature space. The task of predicting plane parameters from a single image is non-trivial, and the complexity is further compounded by multi-task learning. Despite these challenges, our model competes favorably with other, less efficient methods in planar reconstruction that do not offer semantic predictions. To the best of our knowledge, our model also outperforms all other planar reconstruction models in computational efficiency, measured using FPS. It serves as a promising beginning for a real-time multi-task model which can be applied to practical use cases.
Interestingly, the use of multi-view guidance in feature space led to notable cross-task improvements, even with with a lower variety of data used for training. Future research could investigate and quantify the impact of multi-view guidance on models trained with limited data. Exploring other architectural modifications could further enhance model efficiency. Additionally, reconstruction quality can be further enhanced by taking multiple views at inference time. Our work serves as a foundation for advancing semantic plane instance segmentation without sacrificing computational efficiency, striking a balance between efficiency and performance. We hope this work will serve as an inspiration or stepping stone for further research geared towards applications with real-world impact.
Figure 7: 3D visualization of error metrics using predicted parameters. First column uses RGB for projection. Second column shows absolute difference between plane induced depth and ground truth depth. Third column shows cosine distance between the predicted and GT planes.

### v2

@@ 4.1 Setup & Training details
For comparison between different model versions, we train a base model initialized with a pretrained ResNet-50 backbone and employ a data augmentation scheme where each sample has a 15% chance of undergoing one of several augmentations, such as a) jitter of brightness, contrast, hue, saturation, b) Planckian jitter (Zini et al.,, 2023), c) Gaussian noise, or d) motion blur. We use learning rate warm-up for the first 2000 steps starting from a learning rate of $1\mathrm{e}{-6}$ and increases until $2\mathrm{e}{-4}$. After the initial warm-up period, the learning rate is reduced by a factor of 0.1 given no improvement to the validation loss. For quicker and more fair comparison of model variations, a base model with the best validation loss was saved at epoch 9 and used as initialization to our main models, which were trained for 11 additional epochs. We employ early stopping if validation loss fails to improve for 5 consecutive epochs and save the model with best validation performance as well as the last checkpoint. For evaluation, we take the best of either saved model. The additional models trained using the base model initialization do not use data augmentation, and have 500 steps of learning rate warm-up starting from $1\mathrm{e}{-6}$ to $1\mathrm{e}{-5}$. We use a batch size of 32 for the single view model with gradient accumulation to mitigate the higher instability associated with multi-task learning. We train the models on a single NVIDIA Ampere A100 GPU. For evaluation and FPS calculation, we use a single NVIDIA GeForce RTX 3090 GPU for all models.
Table 1: Model comparison results on ScanNet dataset for variations of SOLOP model and other single-image planar reconstruction methods.
@@
PlaneTR 0.178 0.133 0.365 0.215 0.768 0.930 0.977 - - 15
PlaneRCNN 0.165 0.070 0.278 0.187 0.780 0.954 0.991 0.193 - 7
$*{}^{*}$ 0.143 0.059 0.276 0.185 0.813 0.960 0.990 0.416 0.314 38
$*{}^{*}$ 0.134 0.052 0.259 0.178 0.832 0.964 0.991 0.389 0.267 43
$*{}^{*}$ 0.136 0.054 0.261 0.177 0.832 0.962 0.991 0.427 0.344 43
$*{}^{*}$ 0.133 0.052 0.259 0.177 0.833 0.964 0.992 0.434 0.347 43
$*{}^{*}$ = Ours
@@ 4.2 Dataset
For training and evaluation, we use the ScanNet dataset which contains RGB-D images from video sequences, totalling 2.5 million views complete with camera parameters and instance level semantics (Dai et al.,, 2017). The ground truth plane instance annotations for instance masks and plane parameters are generated by the authors of PlaneRCNN, and we follow the same process for filtering and preprocessing the planes (Liu et al.,, 2019). We also obtain the corresponding plane instance semantics from the metadata of the plane annotations. The ground truth plane data often exhibited issues such as over-segmented, rough edges, or missing plane instances, as planes with a depth error above a $0.1$ meter threshold were omitted. For multi-view guidance training, we take sample pairs which are 10 time-steps away. In some cases, a neighbouring ground truth plane image might contain a segment which is missing in the source view, and vice versa. For the single-view model, we use 60,000 random samples from the training set and 10,000 from the validation set. For the multi-view model, we use 30,000 neighboring pairs for training and 5,000 pairs for validation.
Figure 4: Per-pixel recall at varying depth thresholds in meters.
@@ 4.3 Comparison
Our model is most comparable to the PlaneAE model from Yu et al., (2019) and PlaneTR model from Tan et al., (2021), primarily due to the speed of prediction and the fact that they predict plane parameters directly using a single image as input. Given the inconsistent quality of ground truth plane data, the authors of PlaneMVS manually selected stereo pairs for the test set, which contained samples with more complete plane annotations (Liu et al.,, 2022). We run our evaluations using the same test set. For a fair comparison, we train the PlaneAE model for a total of 20 epochs using a ResNet-50 backbone and the same data with an input size of 480 x 640. The original model was trained using an input size of 192 x 256, resulting in a higher FPS. To align with our training regimen, we train PlaneAE for 11 epochs using 60,000 samples and an additional 9 epochs with 100,000 random samples. We retain the original training configuration of the authors (Yu et al.,, 2019). We use the same approach for retraining the PlaneTR model, and generate the required line segments using HAWPv3 (Xue et al.,, 2023). While PlaneRCNN also takes a single image at inference time, its slower inference speed makes it a less direct comparison. We run evaluations on the provided model from authors Liu et al., (2019).
@@ 4.4 Evaluation Metrics
We follow previous methods (Yu et al.,, 2019; Liu et al.,, 2019) and calculate the per-pixel depth recall at varying thresholds in meters, shown in Fig. 4. We also calculate standard depth and detection metrics for a comprehensive evaluation of model performance. Average Precision (AP) is used to assess the quality of the predicted masks, and Mean Average Precision (mAP) takes into account the semantic labels by averaging AP across class categories. For depth metrics, we use Absolute Relative Difference (AbsRel), Squared Relative Difference (SqRel), Root Mean Squared Error (RMSE), log RMSE, and delta accuracy (Eigen et al.,, 2014). We also calculate model efficiency using Frames Per Second (FPS). The results of these evaluations are summarized in Table 1, which shows a marked improvement using our architecture.
@@ 4.5 Results
The task of segmentation becomes more challenging when predicting multiple classes, as overlapping masks from different classes are less likely to be suppressed. The oversegmentation issue appears to be more pronounced in the single view model, whereas multi-view guidance using plane features helped to produce more complete and less over-segmented masks. This improvement is likely attributable to feature sharing and the correlation between ground truth plane instance masks and plane parameters.
@@
Quantitative All results are obtained using the selected test set chosen by the authors of PlaneMVS. The authors Liu et al., (2022) manually selected a higher quality set to evaluate on due to the incomplete and imprecise nature of the ground truth plane annotations. The resulting test set contains 949 image pairs. Our quantitative findings from model comparisons, summarized in Table 1, indicate that our multi-view model variant not only matches the performance of the single-view model in depth metrics, but also shows a significant improvement in detection metrics. This demonstrates the efficacy and improved data efficiency in using multi-view guidance via warping in feature space, at least in the case of using shared features for multitask learning. Since all SOLOP variants use a single image at inference time, the FPS result is the same for the versions of the model using 3 feature levels (SOLOP-SV, SOLOP-MV, SOLOP-MV-gw), but significantly reduced for the version with the original 5 level architecture (SOLOP-5lvls). SOLOP-MV-gw achieves better depth recall comparatively (see Fig. 4), while all SOLOP variants outperform the comparison models on standard metrics.
Qualitative We display different types of visual results from our best model in Figures 1, 5, and 6. In contrast to previous works that predicted a binary plane indication, the incorporation of multi-class semantics introduces an added complexity. The change made to the focal loss for category predictions (see Section 3.2) leads to more confident scoring as well as a potential increase in false positives, which is already exacerbated in the case of multi-class predictions. However, we found that raising the score threshold for the final masks partially mitigated this issue. See Fig. 6 for visual results. The structure of the scene is easier to predict than the exact depth, a challenge presented when using a single image for inference. Sample visualizations of the semantic predictions can be found in Fig. 5. Cases of oversegmentation can occur due to prediction of different classes, or different plane orientation, as each mask represents a planar segment associated with a class label. Overall, our model demonstrates robust performance both visually and quantitatively for the task of planar reconstruction with semantic labels.
@@ 5 DISCUSSION AND FUTURE WORK
In this work, we introduce SOLOPlanes, a real-time semantic planar reconstruction network which shows cross-task improvement when using multi-view guidance in feature space. The task of predicting plane parameters from a single image is non-trivial, and the complexity is further compounded by multi-task learning. Despite these challenges, our model competes favorably with other, less efficient methods in planar reconstruction that do not offer semantic predictions. To the best of our knowledge, our model also outperforms all other planar reconstruction models in computational efficiency, measured using FPS. Our work advances semantic plane instance segmentation without sacrificing computational efficiency, striking a balance between efficiency and performance. We hope it will serve as an inspiration or stepping stone for further research geared towards applications with real-world impact.